United States GDP per person was approximately $59,500 in 2017 [CIA World Factbook 2018]. Measured in imilar dollars, Canadas GDP per person was approximately S48,300 in 2017 . Measured in a] In percentage terms, how much greater w as US GDP per person than Canadian GDP per person in 2017? Hint: employ the following formula: Percentage difference-((US value-Canadian valueycanadian value) x 100% b] Assume US GDP per person grows 2 percent per year over the subsequent 100 years, while Canadian GDP grows 1 per cent per year over the subsequent 100 years. Calculate the size of US GDP per person in 2117.
media%2F94f%2F94f25ecb-8b1c-45db-9a32-cd
Answer #1

a) Percentage difference = (59500 - 48300)/48300 x 100 = 23.19%

b) Size of US GDP per person in 2117 = 59500 x (1 + 2%)100 = $ 431056.4

Size of Canadian GDP per person in 2117 = 48300 x (1 + 1%)100 = $ 130642.5